OpenStudy (anonymous):

[10.02] What are the solutions for x in the proportion 5 times x minus 6 all over x equals 2 times x all over 2.? x = −12 and x = −8 x = 12 and x = 8 x = −2 and x = −3 x = 2 and x = 3

OpenStudy (anonymous):

\[\frac{ 5x-6 }{ x }=\frac{ 2x }{ 2 }\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Know how to solve it??

OpenStudy (anonymous):

Not really I mean I know that you have to substitute the x values for each of them but when I did it none of them worked

OpenStudy (anonymous):

cross multiply it.

OpenStudy (anonymous):

leads to \[10x^2-12=2x^2\] understnad how I got that?

OpenStudy (anonymous):

i mean 10x

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

\[10x-12=2x^2\] becomes \[2x^2-10x-12=0\]

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[2x^2-10x+12=0\] I mean. Someone is distracting me in real life. Sorry

OpenStudy (anonymous):

\[2(x^2-5x+6)=2(x-3)(x-2)=0\] Know how to solve for x?

OpenStudy (anonymous):

I dont understand which one it would be XD Im sorry i'm just really bad at math

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

anything multiplied by a zero equals zero. so set anything equal to zero.

OpenStudy (anonymous):

Wouldnt the answer be x = 2 and x = 3 ?

OpenStudy (anonymous):

Yes, it would. You are right.